# Detecting and Mitigating Insertion Hallucination in Video-to-Audio Generation

### v3

@@ Abstract
Video-to-Audio generation has made remarkable strides in automatically synthesizing sound for video. However, existing evaluation metrics, which focus on semantic and temporal alignment, overlook a critical failure mode: models often generate acoustic events, particularly speech and music, that have no corresponding visual source. We term this phenomenon Insertion Hallucination and identify it as a systemic risk driven by dataset biases, such as the prevalence of off-screen sounds, that remains completely undetected by current metrics. To address this challenge, we first develop a systematic evaluation framework that employs a majority-voting ensemble of multiple audio event detectors. We also introduce two novel metrics to quantify the prevalence and severity of this issue: IH@vid (the fraction of videos with hallucinations) and IH@dur (the fraction of hallucinated duration). Building on this, we propose Posterior Feature Correction, a novel training-free inference-time method that mitigates IH. PFC operates in a two-pass process: it first generates an initial audio output to detect hallucinated segments, and then regenerates the audio after masking the corresponding video features at those timestamps. Experiments on several mainstream V2A benchmarks first reveal that state-of-the-art models suffer from severe IH. In contrast, our PFC method reduces both the prevalence and duration of hallucinations by over 50% on average, without degrading, and in some cases even improving, conventional metrics for audio quality and temporal synchronization. Our work is the first to formally define, systematically measure, and effectively mitigate Insertion Hallucination, paving the way for more reliable and faithful V2A models.
@@ 1 Introduction
Sound design plays a crucial role in enhancing realism and creating an immersive experience in post-production for films, games, animations, and other multimedia content. While silent videos only convey visual information, sound provides richer cues about temporal rhythm, spatial environment, and emotional tone, allowing viewers to better understand and engage with the scene. For instance, suspenseful footsteps in a horror film, distant airplane roars at an airport, or synchronized impact sounds in action games are all indispensable elements. Traditionally, Foley effects are manually recorded, edited, and mixed by sound designers, which is highly specialized, time-consuming, and difficult to scale. This has motivated the development of automatic sound generation systems.
Recently, the task of Video-to-Audio (V2A) generation has attracted increasing attention as a promising solution. V2A models aim to automatically generate sounds that are semantically relevant and temporally aligned with the video content. State-of-the-art models such as MMAudio (Cheng et al., 2025b) and ThinkSound (Liu et al., 2025) have demonstrated impressive results by learning video-audio alignment from large-scale paired datasets. To evaluate their performance, researchers typically adopt metrics such as FD-VGG (Heusel et al., 2017; Hershey et al., 2017), ISC (Salimans et al., 2016), and DeSync (Ruder et al., 2020), which measure the semantic similarity and temporal synchronization between generated and reference audio. These metrics have driven remarkable progress in making V2A models produce the correct category of sounds at the correct moments.
However, these metrics rely on the unverified assumption that generated sounds must exist in the video. In practice, around 50% of VGGSound samples contain off-screen sounds, almost entirely in the Speech and Music categories, while other events remain on-screen (Figure 1; Zverev et al. (2025a)). This bias predisposes models to hallucinate speech or music when visual cues are weak, yet such errors are not captured by current semantic or temporal metrics, leading to misleading assessments of model reliability.
@@
We refer to this overlooked phenomenon as Insertion Hallucination (IH), which denotes the generation of sound events, especially speech and music, that do not exist in the video (Figure 2 illustrates an example where a model outputs speech or music despite their absence in the visual scene). Through empirical studies, we find that representative models including ThinkSound (Liu et al., 2025) and MMAudio (Cheng et al., 2025a) frequently exhibit IH on non-speech and non-music videos from mainstream datasets such as VGGSound (Chen et al., 2020) and the recently released Kling-Audio-Eval (Wang et al., 2025). For example, a model may produce melodic music while the video only depicts sanding, or generate human voices on a vacuum cleaner video. These results suggest that IH is a systematic and widespread risk in V2A generation, yet remains largely undetected by existing metrics and has not been systematically addressed in prior work.
Figure 1: Distribution of on-screen and off-screen sounds in VGGSound. Around 50% of all samples contain off-screen sounds, which are concentrated in the Speech and Music categories, while other events are almost entirely on-screen. This highlights a dataset bias relevant to hallucination analysis.
@@
To systematically investigate IH, we develop an end-to-end evaluation framework. We begin with an automatic detection pipeline that identifies hallucinated speech and music segments by integrating three audio event detectors: inaSpeechSegmenter (Doukhan et al., 2018), YAMNet (Ellis et al., 2019), and PANNs (Kong et al., 2020), and fusing their outputs with majority voting. The pipeline is then validated on a human-annotated set to assess accuracy. Finally, we introduce two metrics: IH@vid (the fraction of videos containing hallucination) and IH@dur (the fraction of hallucinated duration).
@@
Building on this, we propose a novel inference-time correction method called Posterior Feature Correction (PFC). PFC does not require retraining the model. Instead, it runs a two-pass generation process. In the first pass, the model generates audio and we detect hallucinated segments. In the second pass, we mask the video features at those segments by replacing them with empty features, and regenerate the audio. This forces the model to rely on contextual or label information instead of unreliable visual features, preventing it from degenerating into speech or music hallucinations. Experiments show that PFC significantly reduces IH@vid and IH@dur while preserving standard metrics such as FD-VGG, ISC, and DeSync.
@@
  • We propose Posterior Feature Correction (PFC), a training-free inference-time method that significantly reduces IH while maintaining conventional metrics, and demonstrate its effectiveness on multiple V2A benchmarks.
@@ 2.1 Video-to-Audio Generation
Early work on Video-to-Audio generation was dominated by Generative Adversarial Networks (Chatterjee & Cherian, 2020; Pascual et al., 2017; Ferreira et al., 2022), but recent advances have shifted toward models that produce higher-quality audio with stronger audio-visual alignment. A key direction is improving representation alignment: Diff-Foley (Luo et al., 2023) employs contrastive audio-visual pretraining to learn temporally and semantically aligned features that guide a latent diffusion model, achieving substantial gains in synchronization and relevance.
With the rise of more capable generative models, research has expanded toward controllability and practicality. FoleyCrafter (Zhang et al., 2024) adapts a pre-trained text-to-audio model with semantic and temporal controllers, enabling prompt-based control with precise alignment. Data expansion is another path: MMAudio (Cheng et al., 2025a) unifies video-audio and large-scale text-audio data for richer semantics, while MultiFoley (Chen et al., 2025) conditions on both text and audio for flexible user guidance and high-fidelity synchronization.
The frontier is now moving beyond direct mapping to incorporate reasoning. ThinkSound (Liu et al., 2025) introduces a Chain-of-Thought framework in which a multimodal large language model produces interpretable reasoning steps that guide audio generation, transforming the task into a cognitively driven process.
@@ 2.2 Hallucination in Large Language Models
Hallucination is a core challenge for large-scale AI models, with related research expanding from Large Language Models to the multimodal domain. In Large Language Models, researchers mitigate insertion hallucination by introducing external knowledge bases (Lee et al., 2022) or enhancing internal consistency (Mündler et al., 2023). This issue manifests in Vision-Language Models as object hallucination, where a model describes non-existent objects in an image. The academic community has established dedicated evaluation benchmarks such as POPE (Li et al., 2023) and proposed solutions such as Object-Aware Preference Optimization (Chen et al., 2024b; Compagnoni et al., 2025). Recently, the evaluation of hallucination has extended to the audio-visual domain; for instance, AVHBench (Sung-Bin et al., 2024) designs cross-modal understanding tasks to assess whether a model exhibits audio-driven or video-driven hallucinations. However, while these studies prove that hallucination is a common risk in multimodal models, the evaluation endpoint of all existing work, whether for language, vision, or audio-visuals, is exclusively focused on whether the generated textual output contains hallucinations. The phenomenon of the audio itself being the subject of hallucination, such as a Video-to-Audio model generating sound that contradicts the visual scene, remains an unexplored research gap.
@@ 2.3 Off-screen Sound Generation
A notable recent trend in the V2A field is the research on generating off-screen sound. Many researchers have observed that existing video datasets commonly contain off-screen audio events, and they aim to make models learn and align with this characteristic to generate more complete and immersive holistic soundscapes. For instance, VinTAGe (Kushwaha & Tian, 2025) leverages additional information such as text to assist in generating off-screen sounds, while Action2Sound (Chen et al., 2024a) independently models off-screen ambient audio by separating it from foreground sounds. The importance of this trend is also reflected in the evolution of evaluation methods: VGGSounder (Zverev et al., 2025b) was the first to introduce an off-screen sound dimension into its evaluation framework. By comparing model performance with and without visual cues, it revealed a common ”over-reliance on vision” bias in existing models, thereby emphasizing the importance of independent audio understanding capabilities.
@@ 3.2 Multi-Detector Ensemble Framework
Detecting hallucinations reliably requires addressing the limitations of individual audio classifiers. To this end, we design a multi-detector ensemble that combines three complementary detectors: inaSpeechSegmenter (Doukhan et al., 2018), YAMNet (Ellis et al., 2019), and PANNs (Kong et al., 2020).
@@
    Ensemble Fusion. Final predictions are obtained by majority voting across detectors, balancing precision and recall while reducing detector-specific biases:
@@ 3.3 Evaluation Metrics
4 Posterior Feature Correction
Despite recent advances, state-of-the-art V2A models still exhibit Insertion Hallucination when visual inputs provide insufficient cues for reliable generation. This failure mode reflects a systematic reliance on strong dataset priors, in particular speech and music, whenever the visual signal is ambiguous. To address this issue, we propose Posterior Feature Correction (PFC), a training-free inference method that dynamically masks unreliable video features identified through hallucination detection.
@@ 4.1 Method Motivation
We propose Posterior Feature Correction (PFC), which exploits this signal through a two-stage process: first generate audio to identify problematic regions, then regenerate with visual features masked at those locations. By removing unreliable visual cues, we force the model to rely on more conservative generation strategies and stronger contextual information.
This approach is inspired by self-correction mechanisms in other domains (Madaan et al., 2023; Shinn et al., 2023; Huang et al., 2023), but uniquely leverages the temporal structure of audio-visual alignment for targeted feature intervention. An overview of the process is shown in Figure 3.
@@ 4.2 Algorithm Design
PFC operates in two inference passes:
Stage 1: Detection.
@@
Stage 2: Correction.
We construct corrected features $f^{\prime}_{v}$ by replacing the features at hallucinated timestamps with an empty representation $\emptyset_{v}$, a special empty token provided by the pretrained model. This design preserves temporal structure while suppressing misleading cues:
$f^{\prime}_{v}(t)=\begin{cases}\emptyset_{v},&t\in\bigcup_{\tau\in\mathcal{H}}\tau\\ f_{v}(t),&\text{otherwise}.\end{cases}$
The corrected features $f^{\prime}_{v}$ are then fed back to the same model, yielding a revised output $\hat{a}^{\prime}=G(v,f^{\prime}_{v})$.
This two-stage process exploits hallucination locations as uncertainty indicators and intervenes only where necessary. By removing unreliable cues, the model is forced to rely on contextual information and more conservative generation strategies, thereby reducing IH while preserving semantic accuracy and synchronization elsewhere.
Figure 3: Overview of Posterior Feature Correction (PFC). The method first detects hallucination segments (red) in the initial generation, then masks corresponding video features with a learnable empty token to obtain a corrected output (green).
@@ 5 Experiments
This section presents a comprehensive empirical validation of our proposed framework. We first validate our IH detection pipeline on a human-annotated dataset (Section 5.2). We then apply it to assess the prevalence of Insertion Hallucination (IH) in state-of-the-art models and evaluate our Posterior Feature Correction (PFC) method (Section 5.3). Finally, we analyze PFC’s core components via an ablation study (Section 5.4) and compare it against alternative correction methods (Section 5.5).
@@ 5.1 Experimental Setup
We validated our IH detection pipeline using three audio event detectors, inaSpeechSegmenter, YAMNet, and PANNs, on the human-annotated set described in Appendix 7.1, measuring Precision, Recall, $F_{\beta}$-score, and IoU. We then applied the validated metrics to evaluate two representative V2A systems: MMAudio (Cheng et al., 2025a), a multimodal framework with synchronization and flow-matching generation, and ThinkSound (Liu et al., 2025), a reasoning-based model that uses Chain-of-Thought for visual and temporal modeling. Experiments were conducted on three benchmarks: Kling-Audio-Eval (Wang et al., 2025) (20k clips, 1.9k classes, stricter filtering of off-screen sounds), VGGSound (Chen et al., 2020) (200k clips, 310 classes, known off-screen bias), and AVE (Tian et al., 2018) (4,143 clips, 28 classes, frame-level annotations). Evaluation covered hallucination-specific metrics (IH@vid, IH@dur), distributional metrics (FDPANNs, KLPANNs), semantic/quality metrics (ISC, IB-score), and temporal alignment (DeSync).

### v5

@@ Abstract
Video-to-Audio generation has made remarkable strides in automatically synthesizing sound for video. However, existing evaluation metrics, which focus on semantic and temporal alignment, overlook a critical failure mode: models often generate acoustic events, particularly speech and music, that have no corresponding visual source. We term this phenomenon Insertion Hallucination and identify it as a systemic risk driven by dataset biases, such as the prevalence of off-screen sounds, that remains completely undetected by current metrics. To address this challenge, we first develop a systematic evaluation framework that employs a majority-voting ensemble of multiple audio event detectors. We also introduce two novel metrics to quantify the prevalence and severity of this issue: IH@vid (the fraction of videos with hallucinations) and IH@dur (the fraction of hallucinated duration). Building on this, we introduce HALCON to mitigate IH. HALCON follows a three-stage procedure: it first generates initial audio to expose hallucinated segments, then identifies and masks the corresponding unreliable video features, and finally regenerates the audio using the corrected conditioning. Experiments on several mainstream V2A benchmarks first reveal that state-of-the-art models suffer from severe IH. In contrast, our HALCON method reduces both the prevalence and duration of hallucinations by over 50% on average, without degrading, and in some cases even improving, conventional metrics for audio quality and temporal synchronization. Our work is the first to formally define, systematically measure, and effectively mitigate Insertion Hallucination, paving the way for more reliable and faithful V2A models.
@@ 1 Introduction
Sound design is essential for realism and immersion in films, games, animations, and other multimedia content, as audio provides temporal, spatial, and emotional cues that silent visuals alone cannot convey. Traditional Foley production, which involves manually recording, editing, and mixing sound effects, is highly specialized and difficult to scale, motivating the development of automatic sound generation systems. Recent progress in Video-to-Audio (V2A) generation has shown promise in this direction: models such as MMAudio [cheng2025mmaudio] and ThinkSound [liu2025thinksound] learn semantic and temporal audio-video alignment from large-scale paired datasets and achieve impressive performance. Their evaluations typically rely on metrics like FD-VGG [fid, vggish], ISC [isc], and DeSync [desync], which assess semantic similarity and temporal synchronization and have driven substantial advances in generating the right sounds at the right moments.
@@
However, we observe that existing V2A models frequently generate sounds that do not correspond to the visual content—for example, producing melodic music for a sanding scene or synthesizing human voices in a vacuum-cleaner video. As illustrated in Figure 2, such errors often stem from biased audio–visual co-occurrence patterns in training data, which encourage models to associate unrelated events. Yet current evaluation metrics fail to capture these mistakes. In practice, around 50% of VGGSound samples contain off-screen sounds, most of which fall into speech or music categories (Figure 1; vggsounder). This bias predisposes models to hallucinate speech or music when visual cues are weak, while remaining entirely undetected by existing metrics, leading to misleading assessments of model reliability. We refer to this overlooked phenomenon as Insertion Hallucination (IH), which denotes the generation of sound events that have no visual counterpart. Through empirical studies, we find that models such as ThinkSound [liu2025thinksound] and MMAudio [cheng2025mmaudio] frequently exhibit IH on mainstream datasets including VGGSound [vggsound] and Kling-Audio-Eval [klingfoley]. These findings indicate that IH is a systematic and pervasive risk in V2A generation, yet it remains largely undetected by existing evaluation metrics and unaddressed in prior work.
(a) Overview of off-screen sound statistics, including its prevalence and the proportions of speech and music within off-screen events.
(b) Category-level distribution of common off-screen sounds, showing that various forms of speech dominate.
Figure 1: Statistics of off-screen sound in the VGGSound dataset, illustrating its high prevalence and the strong dominance of speech and music within off-screen events.
@@
To systematically investigate IH, we develop an end-to-end evaluation framework. We begin with an automatic detection pipeline that identifies hallucinated segments by integrating three audio event detectors: inaSpeechSegmenter [inaseg], YAMNet [yamnet], and PANNs [panns], and fusing their outputs through majority voting. The pipeline is then validated on a human-annotated set to assess its accuracy. Finally, we introduce two metrics, IH@vid (the fraction of videos containing hallucination) and IH@dur (the fraction of hallucinated duration), to quantify the frequency and severity of IH in V2A models.
@@
Building on this, we propose a novel inference-time conditioning method called HALCON. HALCON does not require retraining the model and instead operates in three stages. In Stage 1, the model generates an initial audio prediction from the input video. In Stage 2, we detect hallucinated segments in the generated audio and correct them by masking the corresponding video features with empty features. In Stage 3, the model regenerates the audio using the corrected video features. This procedure encourages the model to rely on contextual or label information rather than unreliable visual cues, preventing it from degenerating into speech or music hallucinations. Experiments show that HALCON significantly reduces IH@vid and IH@dur while preserving standard metrics.
@@
  • We propose HALCON, a HAlucination-enhanced CONditioning method that significantly reduces IH while maintaining conventional metrics, and demonstrate its effectiveness on multiple V2A benchmarks.
@@ 2.1 Video-to-Audio Generation
Early work on Video-to-Audio generation was dominated by Generative Adversarial Networks [chatterjee2020sound2sight, pascual2017segan, ferreira2022generation], but recent advances have shifted toward models that produce higher-quality audio with stronger audio-visual alignment. A key direction is improving representation alignment: Diff-Foley [luo2023diff] employs contrastive audio-visual pretraining to learn temporally and semantically aligned features that guide a latent diffusion model, achieving substantial gains in synchronization and relevance.
With the rise of more capable generative models, research has expanded toward controllability and practicality. FoleyCrafter [zhang2024foleycrafter] adapts a pre-trained text-to-audio model with semantic and temporal controllers, enabling prompt-based control with precise alignment. Data expansion is another path: MMAudio [cheng2025mmaudio] unifies video-audio and large-scale text-audio data for richer semantics, while MultiFoley [chen2025video] conditions on both text and audio for flexible user guidance and high-fidelity synchronization.
The frontier is now moving beyond direct mapping to incorporate reasoning. ThinkSound [liu2025thinksound] introduces a Chain-of-Thought framework in which a multimodal large language model produces interpretable reasoning steps that guide audio generation, transforming the task into a cognitively driven process.
@@ 2.2 Hallucination in Large Language Models
Hallucination is a core challenge for large-scale AI models, with related research expanding from Large Language Models to the multimodal domain. In Large Language Models, researchers mitigate insertion hallucination by introducing external knowledge bases [lee2022factuality] or enhancing internal consistency [mundler2023self]. This issue manifests in Vision-Language Models as object hallucination, where a model describes non-existent objects in an image. The academic community has established dedicated evaluation benchmarks such as POPE [Li-hallucination-2023] and proposed solutions such as Object-Aware Preference Optimization [chen2024multi, compagnoni2025mitigating]. Recently, the evaluation of hallucination has extended to the audio-visual domain; for instance, AVHBench [sung2024avhbench] designs cross-modal understanding tasks to assess whether a model exhibits audio-driven or video-driven hallucinations. However, while these studies prove that hallucination is a common risk in multimodal models, the evaluation endpoint of all existing work, whether for language, vision, or audio-visuals, is exclusively focused on whether the generated textual output contains hallucinations. The phenomenon of the audio itself being the subject of hallucination, such as a Video-to-Audio model generating sound that contradicts the visual scene, remains an unexplored research gap.
@@ 2.3 Off-screen Sound Generation
A notable recent trend in the V2A field is the research on generating off-screen sound. Many researchers have observed that existing video datasets commonly contain off-screen audio events, and they aim to make models learn and align with this characteristic to generate more complete and immersive holistic soundscapes. For instance, VinTAGe [kushwaha2025vintage] leverages additional information such as text to assist in generating off-screen sounds, while Action2Sound [chen2024action2sound] independently models off-screen ambient audio by separating it from foreground sounds. The importance of this trend is also reflected in the evolution of evaluation methods: VGGSounder [zverev2025vggsounder] was the first to introduce an off-screen sound dimension into its evaluation framework. By comparing model performance with and without visual cues, it revealed a common ”over-reliance on vision” bias in existing models, thereby emphasizing the importance of independent audio understanding capabilities.
@@ 3.2 Multi-Detector Ensemble Framework
Detecting hallucinations reliably requires addressing the limitations of individual audio classifiers. To this end, we design a multi-detector ensemble that combines three complementary detectors: inaSpeechSegmenter [inaseg], YAMNet [yamnet], and PANNs [panns].
@@
    Ensemble Fusion. To combine the outputs of multiple detectors, we consider three fusion strategies: (1) AND, which marks a segment as hallucinated only if all detectors agree; (2) OR, which accepts any detector’s positive prediction; and (3) Majority Vote (MV), which assigns a hallucination label when at least half of the detectors vote positive. The MV rule is formally written as:
@@
    A full comparison of these fusion strategies and the justification for choosing MV are presented in Sec. 5.2.
@@ 3.3 Evaluation Metrics
Figure 3: Overview of our proposed method HALCON.
4 Hallucination-enhanced Conditioning
Despite recent advances, state-of-the-art V2A models still exhibit Insertion Hallucination when visual inputs provide insufficient cues for reliable generation. This failure mode reflects a systematic reliance on strong dataset priors, in particular speech and music, whenever the visual signal is ambiguous. To address this issue, we propose HALCON, a HALlucination-enhanced CONditioning method that dynamically masks unreliable video features identified through hallucination detection.
@@ 4.1 Method Motivation
We propose HALCON, which exploits this signal through a three-stage process: first generate audio to identify problematic regions, then detect and correct the corresponding visual features, and finally regenerate the audio with the corrected features. By removing unreliable visual cues, we force the model to rely on more conservative generation strategies and stronger contextual information.
This approach is inspired by self-correction mechanisms in other domains [madaan2023self, shinn2023reflexion, huang2023large], but uniquely leverages the temporal structure of audio-visual alignment for targeted feature intervention. An overview of the process is shown in Figure 3.
@@ 4.2 Algorithm Design
HALCON operates in three inference stages:
Stage 1: Initial Generation.
Given an input video $v$ with visual features $f_{v}$, the model generates an initial audio prediction $\hat{a}=G(v,f_{v})$.
Stage 2: Hallucination Detection and Feature Correction.
@@
$f^{\prime}_{v}(t)=\begin{cases}\emptyset_{v},&t\in\bigcup_{\tau\in\mathcal{H}}\tau,\\ f_{v}(t),&\text{otherwise}.\end{cases}$
Stage 3: Re-Generation.
The corrected features $f^{\prime}_{v}$ are fed back into the same model to obtain a revised audio output $\hat{a}^{\prime}=G(v,f^{\prime}_{v})$.
This three-stage procedure uses hallucination locations as uncertainty indicators and intervenes only where necessary. By removing unreliable visual cues, the model is encouraged to rely on contextual or label information, thereby reducing IH while preserving semantic accuracy and temporal synchronization elsewhere.
@@ 5 Experiments
This section presents a comprehensive empirical validation of our proposed framework. We first validate our IH detection pipeline on a human-annotated dataset (Section 5.2). We then apply it to assess the prevalence of Insertion Hallucination (IH) in state-of-the-art models and evaluate HALCON (Section 5.3). Finally, we analyze HALCON’s core components via an ablation study (Section 5.5) and compare it against alternative correction methods (Section 5.6).
Method TP (%) FP (%) FN (%) IoU (%)
AND 13.66 1.28 11.39 51.90
OR 21.71 7.49 3.34 66.72
MV 18.46 3.14 6.59 65.47
(a) Percentage-based comparison of TP, FP, FN, and IoU.
(b) $F_{\beta}$ curves across varying $\beta$ values.
Figure 4: Combined evaluation of fusion strategies. The table reports percentage-based metrics, while the curves illustrate the behavior of $F_{\beta}$ under different precision–recall weightings.
@@ 5.1 Experimental Setup
We validated our IH detection pipeline using three audio event detectors, inaSpeechSegmenter, YAMNet, and PANNs, on the human-annotated set described in Appendix, measuring Precision, Recall, $F_{\beta}$-score, and IoU. We then applied the validated metrics to evaluate two representative V2A systems: MMAudio [cheng2025mmaudio], a multimodal framework with synchronization and flow-matching generation, and ThinkSound [liu2025thinksound], a reasoning-based model that uses Chain-of-Thought for visual and temporal modeling. Experiments were conducted on three benchmarks: Kling-Audio-Eval [klingfoley] (20k clips, 1.9k classes, stricter filtering of off-screen sounds), VGGSound [vggsound] (200k clips, 310 classes, known off-screen bias), and AVE [tian2018ave] (4,143 clips, 28 classes, frame-level annotations). Evaluation covered hallucination specific metrics (IH@vid, IH@dur), distributional metrics (FDPANNs, KLPANNs), semantic/quality metrics (ISC, IB-score), and temporal alignment (DeSync).